# Max Getter Chair for Geometry and Analysis

### v1

@@ II-B Energy conservation
The infinite sum $\sum_{\ell=N}^{\infty}\sum_{p\in\mathcal{P}^{(\ell)}}\left\lVert\text{}S[p]f\text{}\right\rVert^{2}$ provides a natural measure to quantify the information contained in the layers of depth $\geq N$. This quantity is bounded above by $W_{N}(f)$. Assuming that $Sf$ is nontrivial, truncating the (possibly infinitely deep) feature extractor after the first $N$ layers looses at most $W_{N}(f)/\left\lVert\text{}Sf\text{}\right\rVert^{2}$ percent of the total energy that is distributed across the entire network, suggesting that a fast decay of this quantity is desirable (see also [7]). Rapid energy decay across layers is also provably closely related to the stability of scattering CNNs with respect to pertubations of the input [5, Corollary 2.15], [14, Theorem 5].
@@ II-C A generic rate for energy decay
While the case $E^{(\ell)}=\emptyset$ may occur in the very general case, we have $E^{(\ell)}\neq\emptyset$ in many practically relevant settings. In fact, the setting is inspired by that of scattering CNNs over general measure spaces [14], where $A^{(\ell)}=H^{t}$ is the diffusion operator defined in [14, Equ. (8),(10)] for some $t>0$, so $E^{(\ell)}=\{(g(\lambda_{k})^{2t},\varphi_{k})\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in I\}$ in the notation of [14].
@@
Let us now continue the discussion from above regarding the concreteness and applicability of our results. First, note that the constraint on the measure space is satisfied for the counting measure, in which case $C_{\mathcal{M}^{(\ell)}}=1$. This may even be regarded as the most practically relevant measure space, having a feature engineer in mind who works with whatever feature extractor on a digital computer. Second, there is a whole class of feature extractors, for which there exists a constant eigenfunction, $\eta\equiv C_{A^{(\ell)}}^{1/2}=\mu^{(\ell)}(M^{(\ell)})^{-1/2}$. Indeed, for scattering CNNs defined via diffusion operators as in [14, Equ. (8)], the eigenfunction $\varphi_{0}$ of $\mathcal{L}$ (in the notation of [14]) is often constant [14, Remark 3], and we have $(1,\varphi_{0})\in E^{(\ell)}$ in the context of the above corollary; see also [14, Theorems 2, 5] and [5, Section 5.1], which are based on similar assumptions.
@@ III-A LCA group scattering
For $g\in L^{1}(G)\cap L^{2}(G)$, let $\mathcal{C}_{g}\in\mathcal{B}(L^{2}(G))$ be the convolution operator $f\mapsto f*g.$ The filters $\Psi\cup\{\chi\}$ induce a scattering CNN [5, 16], with identical architecture for each layer,
@@ III-B A setting-specific rate for energy decay
The next theorem confirms the intuition that the extracted features tend to carry more energy if the frequency gap $\Gamma_{\Psi}$ of the high-pass filters $\Psi$ is rather large. This relates to the fact that the output-generating low-pass filter $\chi$ then captures more information of the processed signal.
@@
$W_{N}(f)\leq\left(1-S^{-1}\right)^{N-1}\cdot\left(\left\lVert\text{}f\text{}\right\rVert^{2}-\left\lVert\text{}f*\chi\text{}\right\rVert^{2}\right).$
@@
The proof of Theorem III.2 makes use of the (accessible) Fourier analysis for abelian (compact) groups, whose analogs are more involving in the nonabelian case. However, we can apply Corollary II.6 if $G$ is nonabelian and finite, which still yields exponential energy decay with a slightly worse basis compared with the abelian case. In fact, taking $(\lambda,\eta)=(1,\mu_{G}(G)^{-1/2}\cdot\mathds{1}_{G})$ in Corollary II.6, we find that $\chi*\chi^{*}*\eta=\eta\geq\mu_{G}(G)^{-1/2}$. Combining with $\mu_{G}(\{e_{G}\})/\mu_{G}(G)=1/\#G$ yields, for all $f\in\mathbb{C}^{G}$, $N\in\mathbb{N}$,
$W_{N}(f)\leq\left(1-1/\#G\right)^{N-1}\cdot\left(\left\lVert\text{}f\text{}\right\rVert^{2}-\left\lVert\text{}f*\chi\text{}\right\rVert^{2}\right).$
@@
$W_{N}(f)\leq\alpha(\Psi,\Gamma)^{N-1}\cdot\left(\left\lVert\text{}f\text{}\right\rVert^{2}-\left\lVert\text{}f*\chi\text{}\right\rVert^{2}\right),$ (9)
@@
Theorem III.4 can be seen to recover [9, Proposition 3.3].
Our results contribute to the understanding of the energy distribution in deep feature extractors and, as a by-product, partially, positively answer the question of the applicability of a stability result for scattering CNNs posed in [14, Remark 7].

### v2

@@ II-B Energy conservation
The infinite sum $\sum_{\ell=N}^{\infty}\sum_{p\in\mathcal{P}^{(\ell)}}\left\lVert\text{}S[p]f\text{}\right\rVert^{2}$ provides a natural measure to quantify the information contained in the layers of depth $\geq N$. This quantity is bounded above by $W_{N}(f)$. Assuming that $Sf$ is nontrivial, truncating the (possibly infinitely deep) feature extractor after the first $N$ layers looses at most $W_{N}(f)/\left\lVert\text{}Sf\text{}\right\rVert^{2}$ percent of the total energy that is distributed across the entire network, suggesting that a fast decay of this quantity is desirable (see also [7]). Rapid energy decay across layers is also provably closely related to the stability of scattering CNNs with respect to pertubations of the input [5, Corollary 2.15], [13, Theorem 5].
@@ II-C A generic rate for energy decay
While the case $E^{(\ell)}=\emptyset$ may occur in the very general case, we have $E^{(\ell)}\neq\emptyset$ in many practically relevant settings. In fact, the setting is inspired by that of scattering CNNs over general measure spaces [13], where $A^{(\ell)}=H^{t}$ is the diffusion operator defined in [13, Equ. (8),(10)] for some $t>0$, so $E^{(\ell)}=\{(g(\lambda_{k})^{2t},\varphi_{k})\leavevmode\nobreak\ |\leavevmode\nobreak\ k\in I\}$ in the notation of [13].
@@
Let us now continue the discussion from above regarding the concreteness and applicability of our results. First, note that the constraint on the measure space is satisfied for the counting measure, in which case $C_{\mathcal{M}^{(\ell)}}=1$. This may even be regarded as the most practically relevant measure space, having a feature engineer in mind who works with whatever feature extractor on a digital computer. Second, there is a whole class of feature extractors, for which there exists a constant eigenfunction, $\eta\equiv C_{A^{(\ell)}}^{1/2}=\mu^{(\ell)}(M^{(\ell)})^{-1/2}$. Indeed, for scattering CNNs defined via diffusion operators as in [13, Equ. (8)], the eigenfunction $\varphi_{0}$ of $\mathcal{L}$ (in the notation of [13]) is often constant [13, Remark 3], and we have $(1,\varphi_{0})\in E^{(\ell)}$ in the context of the above corollary; see also [13, Theorems 2, 5] and [5, Section 5.1], which are based on similar assumptions.
@@ III-A LCA group scattering
For $g\in L^{1}(G)\cap L^{2}(G)$, let $\mathcal{C}_{g}\in\mathcal{B}(L^{2}(G))$ be the convolution operator $f\mapsto f*g.$ The filters $\Psi\cup\{\chi\}$ induce a scattering CNN [5, 16], with each layer sharing the same architecture,
@@ III-B A setting-specific rate for energy decay
The next theorem confirms the intuition that the extracted features tend to carry more energy when the frequency gap $\Gamma_{\Psi}$ is large and the Fourier supports of $\Psi$ are uniformly small. Part of this stems from the low-pass filter $\chi$ capturing more signal information, though the full picture is more nuanced.
@@
$W_{N}(f)\leq\left(\left\lVert\text{}f\text{}\right\rVert^{2}-\left\lVert\text{}f*\chi\text{}\right\rVert^{2}\right)\cdot\left(1-1/S\right)^{N-1}.$
@@
The proof of Theorem III.2 makes use of the (accessible) Fourier analysis for abelian (compact) groups, whose analogs are more involving in the nonabelian case. If $G$ is finite nonabelian, Corollary II.6 still ensures exponential energy decay, albeit with a slightly worse basis compared with the abelian case. In fact, under the low-pass condition $|\int\chi\>\mathrm{d}\mu_{G}|^{2}=1$, taking $(\lambda,\eta)=(1,\mu_{G}(G)^{-1/2}\cdot\mathds{1}_{G})$ in Corollary II.6 yields $\chi*\chi^{*}*\eta=\eta\geq\mu_{G}(G)^{-1/2}$. Since $\mu_{G}(\{e_{G}\})/\mu_{G}(G)=1/\#G$, we have, for all $f\in\mathbb{C}^{G}$, $N\in\mathbb{N}$,
$W_{N}(f)\leq\left(\left\lVert\text{}f\text{}\right\rVert^{2}-\left\lVert\text{}f*\chi\text{}\right\rVert^{2}\right)\cdot\left(1-1/\#G\right)^{N-1}.$
@@
$W_{N}(f)\leq\left(\left\lVert\text{}f\text{}\right\rVert^{2}-\left\lVert\text{}f*\chi\text{}\right\rVert^{2}\right)\cdot\alpha(\Psi,\Gamma)^{N-1},$ (9)
@@
Theorem III.4 generalizes [9, Proposition 3.3] (arbitrary LCA groups, explicit bound for the decay basis).
Our results contribute to the understanding of the energy distribution in deep feature extractors and, as a by-product, partially, positively answer the question of the applicability of a stability result for scattering CNNs posed in [13, Remark 7].